### v1

@@ 2.3 A numerical illustration
We now illustrate the behaviour of all the different SDE samplers discussed previously, in the case where $U(x)$ is given by
@@
corresponding to the case of a truncated Gaussian. For this experiment we set $\epsilon=0.01$. In particular, we study PLMC (5), MYULA (7) as well as two different versions of the MLAm (14) for two different choices of barrier function
$\psi_{1}(x)$ $=-\frac{1}{2}\sum_{i=1}^{2}\left[\log{(x_{i}+1)}+\log{(1-x_{i})}\right],$ (16a)
$\psi_{2}(x)$ $=\frac{1}{2}\psi_{1}(x)+\frac{1}{4}x^{\top}\Sigma^{-1}x.$ (16b)
The choice of $\psi_{1}$ in (16) addresses the constraints associated with $\mathcal{M}$, while in addition $\psi_{2}$ preconditions the dynamics in such a way that the two coordinates move with similar speed.
@@
Figure 1: Histograms for the two different components (top is first component and bottom is the second component) in the case of truncated Gaussian (15). The true density is given by the solid black line.
All these methods lead to some bias with respect to the true invariant distribution that is dictated by the choice of the time-step $\Delta t$. In principle, one can try to eliminate this bias by using a smaller value for it, but this leads to slower mixing and a larger asymptotic variance. When performing our numerical comparisons, we take into account this bias-variance tradeoff to choose the time-step $\Delta t$ for each of the algorithms. For MYULA we use a regularization parameter $\varepsilon=0.01$.
We start our investigation by using $10^{6}$ realisations of each algorithm, with each realisation using $10^{4}$ gradient evaluations. We use the final sample for each of the realisations to produce a histogram for each of the methods. We plot the results in Figure 1. As we can see, when it comes to the $x_{1}$ component both PLMC and MYULA are visually biased, with PLMC putting a lot of probability on the boundary of $\mathcal{M}$, while as expected, MYULA has a large probability of being outside $\mathcal{M}$. On the other hand, MLAm for both choices of barrier function shows no visual bias. No visual bias is observed in the $x_{2}$ component for all the samplers, which is consistent with having no significant probability near the boundary.
To understand the speed of convergence, we calculate the error in Wasserstein $1$-distance between the true distribution and the numerical distribution of each algorithm. Calculating Wassertein errors in dimensions more than one is a highly non-trivial task, but in our case, such a calculation becomes possible by exploiting the independence structure of our probability distribution (more details in the Appendix). We plot in Figure 2 the Wasserstein error in the marginals as well as the total Wasserstein error. We have used 10 copies of the true samples to estimate the Monte Carlo error in the estimator of the $W^{1}$-error, the average error between the true samples is given as the black line while the grey area denotes one standard deviation around the average error. As we can see, consistent with the histograms, the mirror algorithm using $\psi_{1}$ leads to lower bias but the ill-conditioning of $\Sigma^{-1}$ restricts the convergence speed. This is alleviated when the barrier function $\psi_{2}$ is used as the additional preconditioning and leads to faster convergence.
@@ 3.1 Introduction to piecewise deterministic Markov processes
$\frac{\mathrm{d}z_{t}}{\mathrm{d}t}=G(z_{t}),\quad z_{0}=z.$ (17)
The jumps occur at random times $(T_{i})$ that have rate $\lambda(z_{t})$. That is, the inter-jump times $\tau_{i}=T_{i}-T_{i-1}$ have law given by
@@
The framework described above is very general and encompasses a number of important processes. Similarly to the case of SDE based samplers there are a lot of different choices for the triplet $(G,\lambda,Q)$ that would lead to dynamics that are ergodic with respect to (1). We discuss in Section 3.3 some standard choices for which (17) admits an explicit solution.
@@ 3.2 Implementation
An approach to do this is Poisson thinning (see [34]). The idea behind Poisson thinning is that one can simulate an inhomogeneous exponential random variable with rate $m(s)$ as follows: Suppose that there exists a function $M$ such that $m(s)\leq M(s)$ for all $s\geq 0$. Define
@@ 3.3.1 Zig-Zag sampler
To summarise, the dynamics follow a simple linear trajectory $\phi_{t}(x,v)=(x+tv,v)$ between jump times. At jump times one component of the velocity changes sign. Under suitable conditions this process has a unique stationary distribution given by $\rho=\pi\otimes\mathrm{Unif}(\{\pm 1\}^{d})$ (see [11]).
@@
4:  Draw $\mathcal{E}_{i}\sim\mathrm{Exp}(1)$ independently for $i=1,\cdots,d$
@@ 3.3.2 Bouncy particle sampler
$\lambda(x,v)=(v^{\top}\nabla U(x))^{+}+\gamma(x)$ (21)
@@
Here $\nu$ is the velocity marginal of the stationary distribution and is usually taken either to be the normal distribution on $\mathbb{R}^{d}$ or the uniform distribution on the unit sphere of $\mathbb{R}^{d}$. To summarise, the dynamics, similar to the case of the ZZS, follow a simple linear trajectory $\phi_{t}(x,v)=(x+tv,v)$ between jump times. At a jump the velocity is either replaced by a new sample from $\nu$ or is reflected according to the contours of $\nabla U$. Under suitable conditions, this process has a unique stationary distribution given by $\rho(dx,dv)=\pi(dx)\otimes\nu(dv)$ (see [35, 36]).
For implementation, a simple bound for $\lambda$ when $\gamma(x)=\gamma$ is constant and $\nabla U$ is $L$-gradient Lipschitz is given by
@@ 3.4 Subsampling
One of the main advantages of PDMPs over SDE-based approaches is that PDMPs allow for exact subsampling [18] which means that the process can simulate exact samples from $\pi$ while using an unbiased estimator of the gradient $\nabla U$. More precisely, in the context of Bayesian statistics $\pi(x)=p(x|y)$ is the Bayesian posterior density of $x$ given some data $y$. We define $\pi_{0}$ to be the density of the prior, $f(y|x)$ as the likelihood, and $U^{j}(x)=-\log\pi_{0}(x)-K\log f(y^{j}|x)$, with $y=(y^{1},\ldots,y^{K})$ being the data. Then the Bayesian posterior can be written in the form (1) with
@@
In general, there are many cases of practical interest [18, 37] where the function $U$ in (1) is of the form
@@
which only needs evaluation of a single $E_{i}^{j}(x).$ This combined with the properties of Poisson thinning, allows one to implement ZZ-SS without requiring a full pass of the data at each step. In practice, this corresponds to Algorithm 3 which only differs from Algorithm 2 by replacing $\lambda$ with its unbiased estimator in the acceptance rate of the Poisson thinning step.
In principle, when constructing the unbiased estimator one can set $E_{i}^{j}=\partial_{i}U^{j}$, however as shown in [18] it is more efficient to use the idea of control variates. This relies on finding a suitable reference point $\hat{x}$ and then defining
@@
The reference point recommended in [18] is the mode of the distribution, this can be found using a numerical optimisation routine. Although (24) involves the full derivative $\partial_{i}U$, this is only evaluated at the reference point $\hat{x}$ so can be calculated once prior to running the algorithm. A simple bound for $\lambda_{i}^{j}$ when $U$ is $L_{p}$-gradient Lipschitz with respect to the $\lVert\cdot\rVert_{p}$ norm is given by
@@
4:  Draw $\mathcal{E}_{i}\sim\mathrm{Exp}(1)$ independently for $i=1,\cdots,d$
@@ 4 Mirror PDMPs
As we saw in Section 2.3 the mirror map better exploits the geometrical structure of $\mathcal{M}$, and for this reason we will combine ideas of PDMPs and mirror maps. In order to achieve this, we consider ${\zeta}_{t}=\nabla\psi({z}_{t})$ where $\psi$ is a barrier function. Now because of the properties of $\psi$, it is sufficient to ensure that $\zeta_{t}$ has invariant distribution (11). Note that this is an unconstrained distribution and hence one can choose any PDMP which is ergodic with respect to (11) to define $\zeta_{t}$. One can then obtain samples which are ergodic with respect to (1) by applying $\nabla\psi^{*}$ to samples of $\zeta$. We now discuss mirror ZZS which is an implementation based on the ZZS, similarly one can derive the mirror BPS.
@@ 4.1 The Mirror Zig-Zag Sampler
to perform Poisson thinning.
@@
4:  Draw $\mathcal{E}_{i}\sim\mathrm{Exp}(1)$ independently for $i=1,\cdots,d$
@@
Similarly to the discussion on mirror based SDEs it is possible to transform the MZZS into a processes on the primal space. This is itself a PDMP with the following deterministic dynamics
$\begin{cases}\dot{{x}}_{t}=\nabla^{2}\psi({x}_{t})^{-1}{v},\\ \dot{{v}}_{t}=0.\end{cases}$
In addition the rate $\lambda$ satisfies
$\lambda({x},{v})=\widetilde{\lambda}(\nabla\psi({x}),{v})$
while the transition kernel remains unchanged.
Remark 4.2.
@@ 4.1.1 Numerical illustration continued
We now return to the the numerical illustration investigated in Section 2.3 and include the MZZS. As we saw in Section 2.3, the mirror Langevin algorithm with an appropriate barrier function choice exhibits the fastest convergence rate among the algorithms considered and leads to a small bias. Therefore, we will compare this mirror algorithm with MZZS. Since the MLAm is a discrete-time Markov chain, while MZZS gives rise to a continuous-time Markov process, to make this comparison fair, we need to scale time so that the cost of simulating one unit of integration time for both algorithms is roughly equal. We plot the $W^{1}$ error for both algorithms at fixed points of integration time in Figure 3 using $10^{6}$ independent copies each. As we can see, both algorithms converge with the same speed, but notably in the case of the MZZS the method is essentially unbiased as it achieves the same bias as the Monte Carlo error.
@@
For MZZS we have used the mirror map
$\psi_{3}(x)=\frac{1}{2}\psi_{1}(x)+\frac{1}{4}x^{\top}\Sigma^{-\frac{1}{2}}x$
where $\psi_{1}$ is given by (16a). The reasoning for this choice is the same as for (16), except that the relevant map for preconditioning ZZS is different than that for (13). In particular, in the absence of the boundary conditions choosing $\psi(x)=\frac{1}{2}x^{\top}\Sigma^{-1/2}x$ leads to the ZZS for a standard Gaussian in $(\zeta_{t},v_{t}).$
@@ 4.2 Ergodicity of the Mirror Zig-Zag Sampler
for any deterministic initial conditions   $({x},{v})\in\mathcal{M}\times\{-1,1\}^{d}$ and $A\in\mathcal{B}(\mathcal{M}\times\{-1,1\}^{d})$.
Our approach is based on directly employing the results of [11, Theorem 1] for the ZZS $(\zeta_{t},v_{t})$. In particular, if the following assumptions hold then $(\zeta_{t},v_{t})$ is ergodic with respect to $\tilde{\rho}\propto\tilde{\pi}\otimes\mathrm{Unif}(\{\pm 1\}^{d})$, which in turn implies that (28) holds for $(x_{t},v_{t})$.
@@
    Assumption A2. is a technical condition required for the proof of ergodicity in [11].
@@ 5 Numerics and discussions
We now present experimental results demonstrating the performance of the MZZS in a number of different examples. In addition, we compare its performance to state-of-the-art samplers. Throughout all of our experiments unless otherwise stated, we set the parameters for the algorithms as follows: For MLAa we use the step size $1/L_{V}$ with $L_{V}$ being the gradient-Lipschitz constant of $V$, for MLAm we use the step size $1/\tilde{L}$ where $\tilde{L}=L_{V}^{S}+2M_{\psi}L_{V}^{L}$ with $L_{V}^{S}$ being the relative smoothness constant of $U$, $M_{\psi}$ being the self-concordance constant of $\psi$ and $L_{V}^{L}$ being the relative Lipschitz constant of $\nabla U$ (see [38] for details). We implement MZZS using the bound $\tilde{M}$ given by (26), and MBPS using the bound $\tilde{M}$ given by (27). For MBPS the refreshment rate has been manually tuned to obtain the best results. Details on the derivation of the constants will be given in the Appendix.

### v2

@@ 2.3 A numerical illustration
We now illustrate the behaviour of all the different SDE-based samplers discussed previously, in the case where $U(x)$ is given by
@@
corresponding to the case of a truncated Gaussian. For this experiment we set $\epsilon=0.01$. In particular, we study PLMC (5), MYULA (7), and MLAm (14). Initially, we choose our barrier function to be
$\psi_{1}(x)=-\frac{1}{2}\sum_{i=1}^{2}\left[\log{(x_{i}+1)}+\log{(1-x_{i})}\right]$
as it addresses the constraints associated with $\mathcal{M}$.
We start our investigation by using $10^{6}$ realisations of each algorithm, with each realisation using $10^{4}$ gradient evaluations. We use the final sample for each of the realisations to produce a histogram for each of the methods. We plot the results in Figure 1. As we can see, when it comes to the $x_{1}$ component both PLMC and MYULA are visually biased, with PLMC putting a lot of probability on the boundary of $\mathcal{M}$, while as expected, MYULA has a large probability of being outside $\mathcal{M}$. On the other hand, MLAm show no visible bias. In the case of the $x_{2}$ component there is no visible bias for any of the methods.
To understand the speed of convergence, we calculate the error in Wasserstein $1$-distance between the true distribution and the numerical distribution of each algorithm. Calculating Wasserstein errors in dimensions more than one is a highly non-trivial task, but in our case, such a calculation becomes possible by exploiting the independence structure of our probability distribution (more details in the Appendix A). We plot in Figure 2 the Wasserstein error in the marginals as well as the total Wasserstein error. We have used 10 copies of the true samples to estimate the Monte Carlo error in the estimator of the $W^{1}$-error, the average error between the true samples is given as the black line while the grey area denotes one standard deviation around the average error. As we can see, consistent with the histograms, all the three algorithms are biased, with MLAm having the smallest bias. In addition, all the three algorithms suffer from slow convergence due to the ill-conditioning of $\Sigma$.
In order to increase the speed of convergence of the MLAm we now consider the barrier function
$\psi_{2}(x)=\frac{1}{2}\psi_{1}(x)+\frac{1}{4}x^{\top}\Sigma^{-1}x,$
which incorporates the idea of preconditioning, and hence denoted as P-MLAm. Furthermore, to eliminate the bias we consider the Metropolis adjusted algorithm MAMLA. As discussed previously, this is an adjusted discretization of (10) and hence we use a slightly different barrier function given by
$\psi_{3}(x)=\frac{1}{2}\psi_{1}(x)+\frac{1}{4}x^{\top}\Sigma^{-1/2}x.$
For both of the algorithms, there is no visible bias in the histograms, however a closer look of the Wasserstein error shows a small bias for P-MLAm.
@@
(e) MAMLA
Figure 1: Histograms for the two different components for the different algorithms (top first component and bottom second component) in the case of truncated Gaussian (16). The true density is given by the solid black line.
@@ 3.1 Introduction to piecewise deterministic Markov processes
$\frac{dz_{t}}{dt}=G(z_{t}),\quad z_{0}=z.$ (17)
The jumps occur at random times $(T_{i})_{i\geq 1}$ that have rate $\lambda(z_{t})$. That is, the inter-jump times $\tau_{i}=T_{i}-T_{i-1}$ have law given by
@@
The framework described above is very general and encompasses a number of important processes. Similarly to the case of SDE-based samplers, there are a lot of different choices for the triplet $(G,\lambda,Q)$ that would lead to dynamics that are ergodic with respect to (1). We discuss in Section 3.3 some standard choices for which (17) admits an explicit solution.
@@ 3.2 Implementation
An approach to do this is Poisson thinning (see [lewis1979simulation]). The idea behind Poisson thinning is that one can simulate an inhomogeneous exponential random variable with rate $m(s)$ as follows: Suppose that there exists a function $M$ such that $m(s)\leq M(s)$ for all $s\geq 0$. Define
@@ 3.3.1 Zig-Zag sampler
To summarise, the dynamics follow a simple linear trajectory $\phi_{t}(x,v)=(x+tv,v)$ between jump times. At jump times one component of the velocity changes sign. Under suitable conditions this process has a unique stationary distribution given by $\rho=\pi\otimes\mathrm{Unif}(\{\pm 1\}^{d})$ (see [bierkens2019ergodicity]).
@@
4:  Draw $\mathcal{E}_{i}\sim\mathrm{Exp}(1)$ independently for $i=1,\dots,d$
@@ 3.3.2 Bouncy particle sampler
$\lambda(x,v)=(v^{\top}\nabla U(x))^{+}+\gamma(x),$ (21)
@@
Here $\nu$ is the velocity marginal of the stationary distribution and is usually taken either to be the normal distribution on $\mathbb{R}^{d}$ or the uniform distribution on the unit sphere of $\mathbb{R}^{d}$. To summarise, the dynamics, similar to the case of the ZZS, follow a simple linear trajectory $\phi_{t}(x,v)=(x+tv,v)$ between jump times. At a jump the velocity is either replaced by a new sample from $\nu$ or is reflected according to the contours of $\nabla U$. Under suitable conditions, this process has a unique stationary distribution given by $\rho(dx,dv)=\pi(dx)\otimes\nu(dv)$ (see [durmus20, deligiannidis2019exponential]).
For implementation, a simple bound for $\lambda$ when $\gamma(x)=\gamma$ is constant and $U$ is $L$-gradient Lipschitz is given by
@@ 3.4 Subsampling
One of the main advantages of PDMPs over SDE-based approaches is that PDMPs allow for exact subsampling [bierkens2019zig] which means that the process can simulate exact samples from $\pi$ while using an unbiased estimator of the gradient $\nabla U$. More precisely, in the context of Bayesian statistics $\pi(x)=p(x|y)$ is the Bayesian posterior density of $x$ given some data $y$. We define $\pi_{0}$ to be the density of the prior, $f(y|x)$ as the likelihood, and $U^{j}(x)=-\log\pi_{0}(x)-K\log f(y^{j}|x)$, with $y=(y^{1},\ldots,y^{K})$ being the data. Then the Bayesian posterior can be written in the form (1) with
@@
In general, there are many cases of practical interest [bierkens2019zig, WT11] where the function $U$ in (1) is of the form
@@
which only needs evaluation of a single $E_{i}^{j}(x).$ This combined with the properties of Poisson thinning, allows one to implement ZZ-SS without requiring a full pass of the data at each step. In practice, this corresponds to Algorithm 3 which only differs from Algorithm 2 by replacing $\lambda$ with its unbiased estimator in the acceptance rate of the Poisson thinning.
In principle, when constructing the unbiased estimator one can set $E_{i}^{j}=\partial_{i}U^{j}$, however as shown in [bierkens2019zig] it is more efficient to use the idea of control variates. This relies on finding a suitable reference point $\hat{x}$ and then defining
@@
The reference point recommended in [bierkens2019zig] is the mode of the distribution, this can be found using a numerical optimisation routine. Although (24) involves the full derivative $\partial_{i}U$, this is only evaluated at the reference point $\hat{x}$ so can be calculated once prior to running the algorithm. A simple bound for $\lambda_{i}^{j}$ when $U$ is $L_{p}$-gradient Lipschitz with respect to the $\lVert\cdot\rVert_{p}$ norm is given by
@@
4:  Draw $\mathcal{E}_{i}\sim\mathrm{Exp}(1)$ independently for $i=1,\dots,d$
@@ 4 Mirror PDMPs
As we saw in Section 2.3 the mirror map better exploits the geometrical structure of $\mathcal{M}$, and for this reason we will combine ideas of PDMPs and mirror maps. In order to achieve this, we consider ${\zeta}_{t}=\nabla\psi({z}_{t})$ where $\psi$ is a barrier function. Now because of the properties of $\psi$, it is sufficient to ensure that $\zeta_{t}$ is invariant with respect to the push-forward distribution $\tilde{\pi}$ defined through (11). Note that this is an unconstrained distribution and hence one can choose any PDMP which is ergodic with respect to $\tilde{\pi}$ to define $\zeta_{t}$. One can then obtain samples which are ergodic with respect to (1) by applying $\nabla\psi^{*}$ to samples of $\zeta$. We now discuss mirror ZZS which is an implementation based on the ZZS, similarly one can derive the mirror BPS.
@@ 4.1 The Mirror Zig-Zag Sampler
to perform Poisson thinning. Note that, in principle, even though one could derive a general expression for $L_{V}$ in terms of bounds on the derivatives of $U$ and $\psi^{*}$, this expression would be very cumbersome and typically a case by case approach will provide a more useful bound. For all the experiments in this paper we will apply a case by case approach.
Remark 4.1.
Similarly to the discussion on mirror-based SDEs it is possible to transform the MZZS into a processes on the primal space. This is itself a PDMP with the following deterministic dynamics
$\begin{cases}\dot{{x}}_{t}=\nabla^{2}\psi({x}_{t})^{-1}{v_{t}},\\ \dot{{v}}_{t}=0.\end{cases}$
In addition the rate $\lambda$ satisfies
$\lambda({x},{v})=\widetilde{\lambda}(\nabla\psi({x}),{v})$
while the transition kernel remains unchanged.
Observe that these deterministic dynamics can be viewed as a gradient flow in the direction $v$ on the Riemmanian manifold with metric $\nabla^{2}\psi$ which is described in Remark 2.1. Therefore, one can view MZZS as the Riemannian analogue of ZZS where the constant velocity dynamics have become gradient flows with constant direction.
Remark 4.2.
@@
4:  Draw $\mathcal{E}_{i}\sim\mathrm{Exp}(1)$ independently for $i=1,\dots,d$
@@
The proposed algorithm MZZS is similar in spirit to MLAa, see (12), in the sense that both algorithms operate in the dual space using the transformed potential V given by (11) and map back to the primal space using $\nabla\psi^{*}$. In principle, one could construct a mirror PDMP similarly to MLAm by incorporating the geometry of the Riemannian manifold in the dual space. However, doing this is not trivial and the corresponding equations would not typically admit closed form solutions.
@@ 4.1.1 Numerical illustration continued
We now return to the numerical illustration investigated in Section 2.3 and include the MZZS with barrier function $\psi_{3}$. As we saw in Section 2.3, P-MLAm exhibits the fast convergence with small bias, while MAMLA converges with similar speed but with no bias. Therefore, we only compare these mirror algorithms with MZZS. To make this comparison fair, we rescale time so that the cost of simulating one unit of integration time for all those algorithms is roughly equal. We plot the $W^{1}$ error for these algorithms at fixed points of integration time in Figure 3 using $10^{6}$ independent copies. As we can see, all the algorithms converge with the same speed, but notably in the case of the MZZS and MAMLA, the methods are unbiased as they achieve the same bias as the Monte Carlo error.
@@ 4.2 Ergodicity of the Mirror Zig-Zag Sampler
for any deterministic initial conditions $({x},{v})\in\mathcal{M}\times\{-1,1\}^{d}$ and $A\in\mathcal{B}(\mathcal{M}\times\{-1,1\}^{d})$. We will impose the following assumptions:
@@
Under these assumptions we directly employ the results of [bierkens2019ergodicity, Theorem 1] for the ZZS $(\zeta_{t},v_{t})$ to obtain the following theorem.
Theorem 4.1.
If Assumption 4.1 holds then $(\zeta_{t},v_{t})$ is invariant and ergodic with respect to $\tilde{\rho}\propto\tilde{\pi}\otimes\mathrm{Unif}(\{\pm 1\}^{d})$, which in turn implies that (28) holds for $(x_{t},v_{t})$.
@@
    Assumption A2. is a technical condition required for the proof of ergodicity in [bierkens2019ergodicity].
@@ 5 Numerics and discussions
We now present experimental results demonstrating the performance of the MZZS in a number of different examples. In addition, we compare its performance to state-of-the-art samplers. Throughout all of our experiments unless otherwise stated, we set the parameters for the algorithms as follows: For MLAa we use the step size $1/L_{V}$ with $L_{V}$ being the gradient-Lipschitz constant of $V$, for MLAm we use the step size $1/\tilde{L}$ where $\tilde{L}=L_{V}^{S}+2M_{\psi}L_{V}^{L}$ with $L_{V}^{S}$ being the relative smoothness constant of $U$, $M_{\psi}$ being the self-concordance constant of $\psi$ and $L_{V}^{L}$ being the relative Lipschitz constant of $U$ with respect to $\psi$(see [ahn2021efficient] for details). We implement MZZS using the bound $\tilde{M}$ given by (26), and MBPS using the bound $\tilde{M}$ given by (27). For MBPS the refreshment rate has been manually tuned to obtain the best results. For MAMLA, we follow the approach used in the numerical illustration and run the algorithm adaptively, tuning the step size to achieve an acceptance rate of 25%. This value is determined empirically based on the Gamma distribution example. Details on the derivation of the different constants will be given in the Appendix B.